An office furniture supplier makes chairs and desks. Chairs cost $50 each; desks cost $80 each. The business strategy requires that they operate within a $6000 budget. This scenario is best represented mathematically by:

a. 50X1 + 80X2 ≤ 6000

b. Minimize: 6000 + X1 + X2

c. 50X1 + 80X2 ≥ 6000

d. X1 + X2 = 6130
Answer #1

Answer: a: 50x1 + 80x2 <= 6000

Explanation: if no. of chairs= x1 & no. of desks= x2, then

Total cost of chairs= 50*x1 (Chairs cost $50 each;)

Total cost of desks will be= 80*x2 (desks cost $80 each)

And total budget should not exceed 6000: <=6000

Because, Total cost of chairs + Total cost of desks <= Total budget :: 50x1 + 80x2 <= 6000